OpenStudy (anonymous):

Graph the function. Label the vertex and axis of symmetry. y=x^2-8x-20

OpenStudy (anonymous):

Please explain.

OpenStudy (anonymous):

find the factors of 20, that will add up to 8

OpenStudy (anonymous):

so y=(x-10) (x-2)? Then what?

OpenStudy (anonymous):

OK.. check the sign. x-10= 0; x + 2 = 0 ; where y =0

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

x=10 and x =-2

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

the vertex is at the midpoint , so it's between 10 and -2..... imagine a parabola. Mx= -2+10/2

OpenStudy (anonymous):

Mx=8/2=4 so X =4; put in the original formula to find the corresponding My.

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

You need the midpoint. that will be the vertex, (x,y) and the axis of symmetry will run straight down that point.

OpenStudy (anonymous):

Plug in the x=4 to get the y. y=x^2-8x-20

OpenStudy (anonymous):

midpoint means add the two numbers and simply divide by 2, half it.

OpenStudy (anonymous):

But theres two xs, sorry im so confused missed this lesson in class because i was sick

OpenStudy (anonymous):

OK...

OpenStudy (anonymous):

between those two x's is a third x. that is the midpoint you are looking for. Now I'll call it x3; so x3= x1 + x2/ 2

OpenStudy (anonymous):

okay so 6

OpenStudy (anonymous):

no, -2 +10/2= 8/2=4

OpenStudy (anonymous):

the 2 is negative, -2

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now find y

OpenStudy (anonymous):

You'll have 3 points and can now sketch the graph.

OpenStudy (anonymous):

y=-36?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

do you know what the graph looks like?

OpenStudy (anonymous):

no, haha.

OpenStudy (anonymous):